# Differential Contributions of Machine Learning and Statistical Analysis to Language and Cognitive Sciences

### v1

@@ 3.2 Machine learning methods
Machine learning can be categorized mainly into three types: supervised learning, where models predict an outcome based on input data; unsupervised learning, where models identify patterns and relationships in data without any specific outcome to predict; and reinforcement learning, where an agent learns to make decisions by receiving rewards for actions (Bishop and Nasrabadi, , 2006; Hastie et al., , 2009; Murphy, , 2012; Harrington, , 2012; Raschka and Mirjalili, , 2019).
Due to the features of our data, we introduced supervised machine learning models to recognize patterns in the new data. The main purpose of ML is to predict word speech duration. To reduce the training load, we classify the data of word duration into eight ranges. Originally there are 357908 words, and their durations differ from each other, in other words, there are 357908 different values. However, we classified 357908 values into eight ranges. Put it simply, our trained LM models were required to predict which one belongs to a specific range among the eight ones. The data used for training the ML models include age, gender, word length, word frequency. The dataset is divided by 75% as the training one, and the remaining 25% as testing dataset.
@@ 3.2.1 Random forest
Random Forest is an ensemble machine learning algorithm that operates by constructing a multitude of decision trees during training and outputting the class that is the mode of the classes (classification) or mean prediction (regression) of the individual trees. Random Forests correct for decision trees’ habit of overfitting to their training set (Breiman, 2001a, ). During the training phase, random forest creates multiple decision trees. Each tree is trained on a random subset of the data samples, and at each node, a subset of features is randomly chosen to determine the split. For predictions, each tree in the forest votes, and the final class assigned to a sample is based on the majority vote for classification tasks, or an average in case of regression. This aggregation helps to mitigate errors from individual trees and exploit the strength of multiple learners. Random Forest’s performance can typically be assessed using standard metrics such as accuracy for classification tasks. Since the model uses multiple trees, it is usually more robust to overfitting compared to a single decision tree. The final accuracy result is about 51%. The following code is used to implement the task of the classification of different classes on durations.
import pandas as pd
import numpy as np
from sklearn.model_selection import train_test_split
from sklearn.ensemble import RandomForestClassifier
from sklearn.metrics import accuracy_score
import matplotlib.pyplot as plt
from sklearn import tree
@@
# Define the breaks for the ranges of Duration
breaks = np.quantile(df[’Duration’], np.arange(0, 1, 0.2))
labels = [‘1’, ‘2’, ‘3’, ‘4’, ‘5’] # Number of labels matches number of quantile edges minus one
# Cut the ‘Duration’ variable into ranges and assign labels
@@
df[‘range_label’] = pd.cut(df[‘Duration’], bins=breaks, labels=labels, include_lowest=True)
df.dropna(subset=[‘range_label’], inplace=True)
# Feature selection
feature_names = [‘CiteLength’, PhraseLength’, ‘Deletions’, wordlen’, ‘logfreq’]
X = df[feature_names]
y = df[‘range_label’]
# Split the data into training and testing sets
X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.25, random_state=42)
# Create and train the Random Forest Classifier
clf = RandomForestClassifier(n_estimators=100, random_state=42)
clf.fit(X_train, y_train)
# Make predictions
preds_rf = clf.predict(X_test)
accuracy = accuracy_score(y_test, preds_rf)
print(‘‘Accuracy:", accuracy)
#accuracy 51.02712%’
@@
We selected five estimators from 100 ones to visualize them, as shown in Fig. 2. RandomForest model using Python’s matplotlib.pyplot and sklearn.tree libraries. This kind of visualization can be particularly useful for understanding how individual trees in the ensemble make decisions, which can provide insights into the model’s operation.
@@
However, we can streamline the factors in our training dataset. For instance, by selecting only ‘WordLength’ and ‘WordFrequency’, the prediction accuracy achieved was 50.75%. Conversely, using ‘CiteLength’ and ‘PhraseRate’ as factors, the prediction accuracy dropped to 37.13%. Furthermore, combining ‘Deletions’ with ‘WordLength’ yielded a prediction accuracy of 42.36%. By experimenting with various factor combinations, we can deduce the impact of different factors on machine learning accuracy. Evidently, ‘WordLength’ and ‘WordFrequency’ appear to play more significant roles in enhancing machine learning performance.
Figure 2: The five decision trees in the random forest
@@ 3.2.2 Support vector machine
Support vector machine (SVM): SVM is a powerful and versatile supervised machine learning model, particularly well-suited for classification tasks (Hearst et al., , 1998). It works by finding the hyperplane that best divides a dataset into classes with the maximum margin, i.e., the maximum distance between data points of both classes. SVMs are effective in high-dimensional spaces and relatively immune to overfitting, especially in cases where the number of dimensions exceeds the number of samples.
import pandas as pd
import numpy as np
@@
from sklearn.model_selection import train_test_split
from sklearn.svm import SVC
from sklearn.metrics import accuracy_score
# Feature selection
X = df[[‘Age’, ‘Score’, CiteLength’, ‘PhraseLength’, PhraseRate’, ‘Deletions’]]
y = df[‘range_label’]
X = X.dropna()
y = y.loc[X.index]
X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.25, random_state=42)
# Create and train the Support Vector Machine Classifier
clf = SVC(kernel=‘linear’, C=1.0, random_state=42)
clf.fit(X_train, y_train)
preds_svm = clf.predict(X_test)
accuracy = accuracy_score(y_test, preds_svm)
print(‘‘Accuracy:", accuracy)
#accuracy 51.2334%"
@@
We used the SVC from the package “scikit-learn” to create a Support Vector Machine classifier. kernel=‘linear’ specifies that we are using a linear kernel for the SVM. “C=1.0” is the regularization parameter, which controls the trade-off between maximizing the margin and minimizing the classification error. We fit the classifier to the training data using the fit() method. Then, we make predictions on the test data using the predict() method. Finally, we calculate the accuracy of the model using “accuracy_score()” from scikit-learn. We can adjust the kernel type (linear, rbf, poly, etc.) and other hyperparameters of the SVM as needed based on the specific requirements and the characteristics of the dataset. The final accuracy reaches about 51%, which is close to the one predicted by the random forest. The consistent accuracy results show that these factors play stable roles in machine learning. The classifier function from SVM is shown in Fig. 3. We selected four features to visualize how SVM helps classify them.
Figure 3: The two classifiers from the SVM. Each plot selected two factors as x-axis and y-axis.
@@
Following the random forest strategies, we streamlined the factors in our training dataset. For example, using only ‘WordLength’ and ‘WordFrequency,’ we achieved a prediction accuracy of 50.92%. In contrast, CiteLength’ and PhraseRate’ resulted in a lower accuracy of 38.36%, while combining Deletions’ with WordLength’ reached 42.72%. Such results are quite close to those run in the random forest. This approach helped us understand that ‘WordLength’ and ‘WordFrequency’ are more critical for improving performance. In the following statistical analysis, we should treat both factors as control predictors.
3.3 Statistical analysis
@@
Statistical analysis is broadly divided into two main types: descriptive statistics and inferential statistics. Descriptive statistics summarize and describe the features of a dataset through metrics like mean, median, mode, and standard deviation, helping to visualize and understand data distributions. Inferential statistics, on the other hand, use samples of data to make generalizations or predictions about a larger population, employing techniques such as hypothesis testing, regression analysis, and confidence intervals. This distinction allows statisticians and researchers to both understand the data they have and to infer properties about data they do not have, supporting decision-making across fields like economics, medicine, engineering, and social sciences.
@@
We first employed descriptive correlation to explore in the data, and then applied some advanced regression models to make further explorations. Machine learning needs to use training data to train models and do tests in the test data. In contrast, statistical analysis does not require to do this.

### v2

@@ 3.2 Study 1: Machine learning methods
ML can be classified in various ways based on different criteria. One common categorization divides ML into three main types: supervised learning, where models predict an outcome based on input data; unsupervised learning, where models identify patterns and relationships in data without any specific outcome to predict; and reinforcement learning, where an agent learns to make decisions by receiving rewards for actions (Bishop and Nasrabadi, , 2006; Hastie et al., , 2009; Murphy, , 2012; Harrington, , 2012; Raschka and Mirjalili, , 2019). The general mathematical equation for machine learning could be formulated as follows:
$y=f(\mathbf{X})+\varepsilon$ (1)
Where $y$ is the target variable (output), $\mathbf{X}$ is the input variable or feature vector $\mathbf{X}=[x_{1},x_{2},\ldots,x_{n}]$, $f$ is the function we are trying to learn or approximate, $\varepsilon$ represents the error term or noise. Basically, the goal of machine learning is to optimize the parameters in Equation (1). The key to optimize the parameters is to reduce the error or loss. From this perspective, the main purpose of training a machine learning model is to minimize its loss function. The loss function $L$ can be expressed as:
$L=\sum_{i=1}^{n}(y_{i}-f(\mathbf{X}_{i}))^{2}$ (2)
ML can accomplish three primary tasks: clustering, classification, and regression. Various typical machine learning models exist, including linear regression, decision trees, random forests, support vector machines (SVMs), K-nearest neighbors (KNN), Naive Bayes, K-Means, and principal component analysis (PCA). Each model has its unique strengths that make it particularly suited for specific tasks. For instance, K-Means excels at clustering, while random forests and decision trees are adept at classification. Linear regression, on the other hand, is particularly effective for regression tasks.
When discussing specific models, their functions can indeed be more specialized compared to their general purpose. For linear regression exampled in the section 2.3, $f(\mathbf{X})$ might take the form:
$f(\mathbf{X})=w_{0}+w_{1}x_{1}+w_{2}x_{2}+\ldots+w_{n}x_{n}$ (3)
After grasping the fundamental principles of machine learning, we can outline the basic steps for implementing machine learning to perform classification or regression tasks. These steps can be summarized in Table 3.
Table 3: Steps for implementing ML models
Step Implementation Description
1 Define problem and gather Data Clearly state the problem(s) you’re trying to solve and collect relevant, high-quality data.
2 Preprocess data and select features Clean the data, handle missing values, and understand patterns and relationships. Select useful features from the data.
3 Choose model Select an appropriate machine learning algorithm based on the problem type (e.g., classification, clustering, regression).
4 Split data Divide your dataset into training and testing sets.
5 Train model Use the training data to teach the chosen algorithms to make predictions or decisions.
6 Evaluate model Test the model’s performance using the testing data and appropriate metrics (e.g., Precision, Recall, F1-score, Mean Absolute Error etc. )
7 Tune and optimize Adjust (hyper)parameters and refine the model using some strategies (e.g., bootstrap, normalizing the data etc.) to improve its performance.
8 Deploy and Monitor (optional) Integrate the model into the application or system, and continuously monitor its performance in real-world conditions.
Following the steps outlined in Table 3, we can detail the implementation of a machine learning task using the Buckeye corpus. Initially, we define the problem. Machine learning will be employed for clustering, classification, or regression tasks. Given the characteristics of the Buckeye corpus, our goal is to employ supervised machine learning models to identify patterns in new data. The primary objective of this ML task is to predict the ranges of word speech duration, which is a typical classification problem. To alleviate the training burden, we categorize word durations into eight distinct ranges. Originally, there are 357,908 words, each with a unique duration. However, we have grouped these 357,908 unique values into eight categories. Essentially, using other features in the dataset, the chosen ML models are intended to predict which of the eight categories a word’s duration falls into.
Second, referring to Step 3 in Table 3, we could introduce several machine learning models to accomplish the classification task. A variety of models could be selected to implement this classification. Based on the nature of the task and the dataset, we have chosen the Random Forest (RF) and support vector machine (SVM) models to perform the classification. The reason for this is that both RF and SVM perform well in classification tasks, making them typical and reliable ML models for this purpose. Another reason is the specific characteristics of the Buckeye Corpus dataset.
Third, to train an effective ML model, we must carefully select relevant features from the dataset. For this study, we believe that the factors from the Buckeye corpus such as age, gender, word length, word frequency, and speaking rate, among others, could be particularly valuable. Once we have identified these key features, we’ll divide our dataset into two portions: 75% for training and 25% for testing. The larger training set will be used to train our ML model, while the smaller testing set will be reserved for evaluating the model’s performance on unseen data. This approach ensures an unbiased assessment of the model’s effectiveness and generalizability.
Lastly, in most cases, we need to implement strategies to optimize the model to enhance the evaluation results. Probably we could decrease or increase some given features in the model to improve the model’s performance. In most cases, some strategies for each speficic models could be considered. This step can help minimize the loss function described in Equation (2), thereby optimizing the parameters in Equation (1) for improved outcomes.
The following sections demonstrate the implementation of RF and SVM models, illustrating how to implement these steps in practice. We also provide the corresponding Python scripts to execute these steps.
@@ 3.2.1 Random forest
Random Forest is a sophisticated ensemble machine learning algorithm that constructs multiple decision trees during the training phase, addressing the overfitting tendencies often associated with individual decision trees (Breiman, 2001a, ). This method operates by training each tree on a random subset of the data and selecting a random subset of features at each node to determine the optimal split. For predictions, Random Forest employs a democratic approach: in classification tasks, each tree votes, and the majority determines the final class assignment, while in regression tasks, the algorithm calculates the mean of individual tree predictions. This aggregation methodology serves to mitigate errors from individual trees and leverage the collective strength of multiple learners, resulting in enhanced robustness against overfitting. The algorithm’s performance is typically evaluated using standard metrics such as accuracy for classification tasks. This versatile approach has found wide application across various domains, including finance, healthcare, and environmental sciences, due to its ease of use, ability to handle both categorical and continuous variables, built-in feature importance ranking, and efficiency in processing large datasets.
The subsequent Python script demonstrates the implementation of a classification task for different duration classes. This code exemplifies the specific steps required to develop and deploy a machine learning model for the classification purpose designed for predicting the ranges of word speech duration in the Buckeye corpus, as outlined in Table 3. The script serves as a practical illustration of the theoretical framework discussed earlier, showcasing how each step in the machine learning process is translated into executable code.
Listing 1: Python code for loading data and training a RandomForest classifier
1 import pandas as pd
2 import numpy as np
3 from sklearn.model_selection import train_test_split
4 from sklearn.ensemble import RandomForestClassifier
5 from sklearn.metrics import accuracy_score
6 import matplotlib.pyplot as plt
7 from sklearn import tree
8
9 # Load the data
@@
11
12 # Define the breaks for the ranges of ’Duration’
13 breaks = np.quantile(df[’Duration’], np.arange(0, 1, 0.2))
14
15 # Define labels
16 labels = [’1’, ’2’, ’3’, ’4’, ’5’] # Number of labels matches number of quantile edges minus one
17
18 # Cut the ’Duration’ variable into ranges and assign labels
@@
19 df[’range_label’] = pd.cut(df[’Duration’], bins=breaks, labels=labels, include_lowest=True)
20
21 # Drop rows with missing values in ’range_label’
22 df.dropna(subset=[’range_label’], inplace=True)
23
24 # Feature selection
25 feature_names = [’CiteLength’, ’PhraseLength’, ’Deletions’, ’WordLength’, ’LogWordFreq’]
26 X = df[feature_names]
27 y = df[’range_label’]
28
29 # Split the data into training and testing sets
30 X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.25, random_state=42)
31
32 # Create and train the Random Forest Classifier
33 clf = RandomForestClassifier(n_estimators=100, random_state=42)
34 clf.fit(X_train, y_train)
35
36 # Make predictions
37 preds_rf = clf.predict(X_test)
38 accuracy = accuracy_score(y_test, preds_rf)
39 print("Accuracy:", accuracy)
40 # Accuracy: 51.02712%
@@ 26 X = df[feature_names]
The RF model using Python’s matplotlib.pyplot and sklearn.tree libraries. This kind of visualization can be particularly useful for understanding how individual trees in the ensemble make decisions, which can provide insights into the model’s operation.
@@
However, we can streamline the factors (features) in the training dataset. For instance, by selecting only ‘WordLength’ and ‘WordFrequency’, the prediction accuracy achieved was 50.75%. Conversely, using ‘CiteLength’ and ‘PhraseRate’ as factors, the prediction accuracy dropped to 37.13%. Furthermore, combining ‘Deletions’ with ‘WordLength’ yielded a prediction accuracy of 42.36%. By experimenting with various factor combinations, we can deduce the impact of different factors on machine learning accuracy. Evidently, ‘WordLength’ and ‘WordFrequency’ appear to play more significant roles in enhancing machine learning performance.
To enhance the performance of our ML model, we explored the impact of incorporating additional features. We expanded our feature set to include ‘Age’: feature_names = [‘Age’, ‘CiteLength’, ‘PhraseLength’, ‘Deletions’, ‘WordLength’, ‘LogWordFreq’]. Following the same training procedure with this expanded feature set, we observed an improvement in the accuracy score from 51% to 53.34%. Further, when we included ‘Speaker’ as an additional feature, the accuracy score increased to 59.27%. These results show the critical role that feature selection plays in optimizing ML model performance. By carefully choosing and incorporating relevant features, we can significantly enhance the performance of the trained ML models. Specifically, the factor ‘Speaker’ is a signficant feature in training ML classifier.
Despite of some progress in classification, we still need to introduce some optimized strategies to improve the model performance, that is, we could improve the accuracy score, and this step is “Tune and optimize” in Table 3. These strategies for optimizing the RF model include the following. First, feature scaling is introduced to normalize the features and it may improve the model’s accuracy. Second, the implementation of hyperparameter tuning through the function of “GridSearchCV” allows for the optimization of various parameters in this Random Forest Classifier, ensuring that the model is fine-tuned for better predictive performance. Third, we employ feature selection with the function of “SelectFromModel”, identifying and retaining only the most important features based on their importance scores, which can lead to a more efficient and effective model.
Furthermore, the evaluation of model performance could be more comprehensive. To this end, we include a classification report that details precision, recall, and F1-score for each class, providing deeper insights into how well the model performs across different categories. To aid in understanding feature significance, a bar plot visualizing feature importances is also included. Lastly, the breaks calculation for creating duration ranges has been slightly adjusted to ensure that the upper bound is included. Overall, these changes reflect best practices in machine learning and are likely to yield improved accuracy and interpretability in the model’s results. The implementation script is seen below.
1 import pandas as pd
2 import numpy as np
3 from sklearn.model_selection import train_test_split, GridSearchCV
4 from sklearn.ensemble import RandomForestClassifier
5 from sklearn.preprocessing import StandardScaler
6 from sklearn.metrics import accuracy_score, classification_report
7 import matplotlib.pyplot as plt
8 from sklearn.feature_selection import SelectFromModel
9
10 ##the same code
11
12 # Scale the features
13 scaler = StandardScaler()
14 X_train_scaled = scaler.fit_transform(X_train)
15 X_test_scaled = scaler.transform(X_test)
16
17 # Create and train the Random Forest Classifier with hyperparameter tuning
18 param_grid = {
19 ’n_estimators’: [100, 200, 300],
20 ’max_depth’: [None, 10, 20, 30],
21 ’min_samples_split’: [2, 5, 10],
22 ’min_samples_leaf’: [1, 2, 4]
23 }
24
25 rf = RandomForestClassifier(random_state=42)
26 grid_search = GridSearchCV(estimator=rf, param_grid=param_grid, cv=5, n_jobs=-1, verbose=2)
27 grid_search.fit(X_train_scaled, y_train)
28
29 # Get the best model
30 best_rf = grid_search.best_estimator_
31
32 # Feature importance-based selection
33 selector = SelectFromModel(best_rf, prefit=True)
34 X_train_selected = selector.transform(X_train_scaled)
35 X_test_selected = selector.transform(X_test_scaled)
36
37 # Train the model with selected features
38 best_rf.fit(X_train_selected, y_train)
39
40 # Make predictions
41 preds_rf = best_rf.predict(X_test_selected)
42 accuracy = accuracy_score(y_test, preds_rf)
43 print("Accuracy:", accuracy)
44
45 # Print classification report
[…]
47 print(classification_report(y_test, preds_rf))
48
49 # Plot feature importances
50 importances = best_rf.feature_importances_
51 indices = np.argsort(importances)[::-1]
52 selected_features = selector.get_support()
53
54 plt.figure(figsize=(10,6))
55 plt.title("FeatureImportances")
56 plt.bar(range(X_train_selected.shape[1]), importances[indices][:X_train_selected.shape[1]])
57 plt.xticks(range(X_train_selected.shape[1]), [feature_names[i] for i in indices if selected_features[i]], rotation=45)
58 plt.tight_layout()
59 plt.show()
Upon execution of the revised script, we observed a substantial improvement in the model’s performance. The accuracy score increased from the earlier 59.27% to 73.32%, representing a significant enhancement of 14 percentage points. This marked improvement demonstrates the efficacy of the implemented strategies in optimizing the classifier’s performance. These results not only validate our approach but also highlight the potential for further optimization in similar classification problem. Certainly other strategies could be employed to continue optimizing the performance of this classifier. After the classifier meets our expected result, the classifier could be saved and be deployed to fulfill the automatic classification task of word duration for a new dataset of spontaneous speech given the features such as word length, word frequency, PhraseLength are taken.
@@ 3.2.2 Support vector machine
SVM is a powerful and versatile supervised machine learning model, particularly well-suited for classification tasks (Hearst et al., , 1998). It works by finding the hyperplane that best divides a dataset into classes with the maximum margin, i.e., the maximum distance between data points of both classes. SVMs are effective in high-dimensional spaces and relatively immune to overfitting, especially in cases where the number of dimensions exceeds the number of samples.
The following Python script illustrates the implementation of a SVM for classifying word speech durations into different categories. This code exemplifies the step-by-step process of developing and deploying a classification model, specifically tailored to predict duration ranges of words in the Buckeye corpus. Each stage of the implementation corresponds to the machine learning steps outlined in Table 3.
1 import pandas as pd
2 import numpy as np
@@
3 from sklearn.model_selection import train_test_split
4 from sklearn.svm import SVC
5 from sklearn.metrics import accuracy_score
6
7 # Feature selection
8 X = df[[’CiteLength’, ’PhraseLength’, ’PhraseRate’, ’Deletions’, ’WordLength’, LogWordFrequency’]]
9␣␣␣␣y=df[‘range_label’]
10 X = X.dropna()
11 y = y.loc[X.index]
12
13 X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.25, random_state=42)
14
15 # Create and train the Support Vector Machine Classifier
16 clf = SVC(kernel=‘linear’,C=1.0,random_state=42)
17␣␣␣␣clf.fit(X_train,y_train)
18
19␣␣␣␣preds_svm=clf.predict(X_test)
20␣␣␣␣accuracy=accuracy_score(y_test,preds_svm)
21␣␣␣␣print(‘‘Accuracy:",accuracy)
22␣␣␣␣#accuracy51.2334%’
@@ 8 X = df[[’CiteLength’, ’PhraseLength’, ’PhraseRate’, ’Deletions’, ’WordLength’, LogWordFrequency’]]
We used the SVC from the package “scikit-learn” to create a classifier. kernel=‘linear’ specifies that we are using a linear kernel for the SVM. “C=1.0” is the regularization parameter, which controls the trade-off between maximizing the margin and minimizing the classification error. We fit the classifier to the training data using the fit() method. Then, we make predictions on the test data using the predict() method. Finally, we calculate the accuracy of the model using “accuracy_score()” from scikit-learn. We can adjust the kernel type (linear, rbf, poly, etc.) and other hyperparameters of the SVM as needed based on the specific requirements and the characteristics of the dataset. The final accuracy reaches about 51%, which is close to the one predicted by the random forest. The consistent accuracy results show that these factors play stable roles in machine learning. We selected four features to visualize how SVM helps classify them.
@@
Following the RF analysis, we streamlined the factors in our training dataset. For example, using only ‘WordLength’ and ‘WordFrequency,’ we achieved a prediction accuracy of 50.92%. In contrast, CiteLength’ and PhraseRate’ resulted in a lower accuracy of 38.36%, while combining ‘Deletions’ with ‘WordLength’ reached 42.72%. Such results are quite close to those run in the random forest. This approach helps us understand that ‘WordLength’ and ‘WordFrequency’ are more critical for improving performance. In the following statistical analysis, we should treat both factors as control predictors. The addition of the features, such as ‘Age’ and ‘Speaker’, improved the accuracy scores, and the scores are almost similar to the ones in the RF experiments (i.e., 53.37%, 59.42%).
Despite the initial implementation, optimizing the SVM classifier is a critical step in improving the performance of SVM classifier. Several strategies can be employed to enhance the classifier’s accuracy and efficiency. Following the optimization of the SVM classifier, its accuracy score improved to 72.33%, which is comparable to the performance of the optimized Random Forest classifier. The specific strategies and their corresponding implementation script are detailed in Appendix B. These findings highlight that optimizing machine learning implementation is crucial for enhancing model performance. An optimized ML model can then be deployed to process new datasets to make predictions effectively.
3.3 Study 2: Statistical analysis
@@
Statistical analysis is broadly divided into two main types: descriptive statistics and inferential statistics. Descriptive statistics summarizes and describes the features of a dataset through metrics like mean, median, and standard deviation, etc., helping to visualize and understand data distributions. Inferential statistics, on the other hand, use samples of data to make generalizations or predictions about a larger population, employing techniques such as hypothesis testing, regression analysis, and confidence intervals. This distinction allows statisticians and researchers to both understand the data they have and to infer properties about data they do not have, supporting decision-making across fields like economics, medicine, engineering, and social sciences.
@@
We first employed descriptive correlation to explore in the data of the Buckeye corpus, and then applied some advanced regression models to make further explorations. Implementing advanced regression models for statistical analysis is significantly distinct from the implementation of ML models.